# LLINBO: Trustworthy LLM-in-the-Loop Bayesian Optimization

### v1

@@ Abstract
Figure 1: Diagrams of existing methods, which are BO and LLM-assisted BO, and the three proposed algorithms: LLINBO-Transient, LLINBO-Justify and LLINBO-Constrained, introduced in Sec. 2.3, 2.4 and 2.5 respectively.
@@ 1 Introduction
BO has emerged as a powerful tool for black-box optimization, providing a principled framework for balancing exploration and exploitation. BO is particularly useful in scenarios where function evaluations are costly, such as in drug discovery [20], materials science [12, 35], interaction design [23], and hyperparameter tuning [7, 36].
Starting with an initial dataset, BO employs a surrogate model, most commonly a $\mathcal{GP}$. The $\mathcal{GP}$ is capable of quantifying uncertainty and is used to approximate both the mean and variance of the black-box function. The next query point, hereafter referred to as a design, is then selected by maximizing an acquisition function (AF) that quantifies the potential benefit of evaluating a particular point, thereby strategically balancing exploration and exploitation. BO then augments the dataset with the new design–outcome tuple and proceeds sequentially. The past decade has witnessed many success stories for BO, and its theoretical guarantees have been well established for a range of commonly used AFs [32, 17]. These guarantees are typically regret-based, ensuring that, with high probability, one can asymptotically recover an optimal design.
Recently, the few-shot learning capabilities of LLMs and their ability to generate high-quality outputs from minimal examples have made them attractive tools for optimization tasks [38]. In particular, LLMs have shown strong empirical performance over random search [25], largely due to their ability to leverage problem context to fast-track the exploration of promising designs. Intuitively, LLMs act like domain experts, using contextual cues to identify high-quality designs early in the optimization process. At each iteration, different phases of BO, including initial data generation, proposing new designs, and surrogate modeling, are carried out by the LLM through appropriately tailored prompts [25, 38]. These prompts incorporate the current dataset, typically presented as a list of design-response pairs, together with the problem context, enabling the LLM to function as an optimizer. This prompting framework allows LLMs to act as potential agents for black-box optimization without the need for explicit surrogate modeling or large amounts of observed data. We refer to this class of approaches, where LLMs are solely responsible for proposing design candidates and serve as the surrogate model in BO, as LLM-assisted BO.
@@
While very recent LLM-assisted BO [25, 14, 31, 38] approaches have shown promise in suggesting reasonable query designs, several limitations hinder their broader applicability. Most importantly, such approaches lack explicit surrogate modeling and calibrated uncertainty [30], which are critical for managing the exploration–exploitation trade-off [4]. Moreover, LLMs remain inherently opaque, making the aforementioned trade-off difficult to interpret or control. This structural opacity, combined with their inability to quantify uncertainty in a principled way, introduces significant risks, particularly in applications where cost or safety is critical, ultimately undermining theoretical tractability and reliability. For instance, in the case of smooth functions, the predictive capability of $\mathcal{GP}$s, in terms of both the predicted mean and variance as measured by generalization bounds, has a known rate of improvement as more data is gathered [32, 29]. The same result is hard to characterize for LLMs, whose internal mechanisms for interpolating black-box functions are not fully understood and which currently lack calibrated uncertainty estimates.
With this in mind, we propose LLINBO, a framework that combines the contextual reasoning strengths of LLMs with the principled uncertainty quantification offered by statistical surrogates to enable more trustworthy and tractable optimization. To the best of our knowledge, this is the first principled hybrid framework for BO that systematically integrates LLMs and statistical surrogates. To operationalize this collaboration, we introduce a general framework grounded in the philosophy of using LLM-suggested designs to sequentially refine and tailor BO. Within this framework, we propose three approaches, which are inspired by recent developments in collaborative learning, and analyze the theoretical properties of each. Through extensive simulations and a real-world proof-of-concept in 3D printing, we demonstrate the effectiveness and robustness of the proposed methods.
@@
LLMs’ ability to utilize problem context has been actively investigated, with Xie et al. [37] interpreting in-context learning as a form of implicit Bayesian inference. Recent work has also demonstrated that LLMs can generalize effectively from limited in-context information [22, 5], making them particularly promising for black-box optimization, where the objective function is unknown and historical observations are limited [25]. The use of LLMs for optimization is a growing research direction, yet still in its infancy. For example, Liu et al. [24] employed LLMs to solve multi-objective optimization problems, while Guo et al. [14] extended the use of LLMs to a broader set of tasks, including combinatorial optimization. Very recently, Song et al. [31] explored how LLMs can enhance black-box optimization by leveraging textual knowledge and sequence modeling to improve generalization.
While existing efforts have primarily focused on prompting-based strategies for optimization [25, 14, 31, 38, 24], this work specifically targets black-box scenarios and enhances trustworthiness by integrating principled surrogates; thereby addressing the risks of relying solely on LLMs, whose surrogate modeling, uncertainty quantification, and exploration–exploitation behavior remain opaque. At the heart of our approach is the collaboration between an LLM and a statistical surrogate, e.g., a $\mathcal{GP}$. Accordingly, some of the tools developed below draw on principles from collaborative and federated learning [19]. Needless to say, these are extensive fields in their own right, so we only highlight works that are immediately relevant to our paper, in the context of BO. Yue et al. [39] developed a consensus framework for collaborative BO, where the next design to query is selected as a weighted combination, dictated by a dynamically coupled stochastic consensus matrix, of the AF maximizers from all clients in the system, including each client’s own. In [9, 10], Federated Thompson Sampling for BO was proposed, where clients share $\mathcal{GP}$ Random Fourier Features [27]. Each client then samples the next design to query either based on its own features or on those of another randomly selected client. Alternatively, Chen et al. [6] proposed a constraint-sharing approach, where clients resample their surrogates based on shared constraints to determine the next design to evaluate. For a recent overview of collaborative BO, refer to [11, 1]. While this work differs in its goal, its principles have inspired our hybrid collaboration between LLMs and statistical surrogates.
@@ 2.1 Preliminaries
BO aims to find an optimal design that maximizes a black-box function $f:\mathcal{X}\rightarrow\mathbb{R}$ over a domain $\mathcal{X}$, i.e., $x^{\ast}=\arg\max_{x\in\mathcal{X}}f(x)$, by sequentially selecting query designs. Given a total budget of $T$ evaluations, the data at iteration $t\in[T]$ is denoted as $\mathcal{D}_{t-1}=\{(x_{i},y_{i})\}_{i=1}^{t-1}$, where $y_{i}=f(x_{i})+\epsilon_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,\lambda^{2})$.
At time $t$, BO selects the next design to observe by maximizing an AF, $\alpha(x,F_{t-1})$, where $F_{t-1}$ is the posterior belief of $f$ conditioned on $\mathcal{D}_{t-1}$. Specifically, the next design is chosen as
$x_{t}=\mathop{\arg\max}_{x\in\mathcal{X}}\alpha(x,F_{t-1}).$ (1)
After selecting $x_{t}$, a noisy observation $y_{t}=f(x_{t})+\epsilon_{t}$ is obtained, and the dataset is updated as $\mathcal{D}_{t}=\mathcal{D}_{t-1}\cup\{(x_{t},y_{t})\}$. This process is then repeated until $T$ is exhausted. The posterior belief is typically modeled using a $\mathcal{GP}$ [21], which requires a prior mean function $\mu(x)$ (often set to zero) and a kernel function $k(x,x^{\prime})$ encoding the smoothness of the function. This yields a posterior predictive distribution for $f$ given as
@@
with
$\mu_{t-1}(x)$ $=k_{t-1}(x)^{\top}(K+\lambda^{2}I)^{-1}y,$
$\sigma^{2}_{t-1}(x)$ $=k(x,x)-k_{t-1}(x)^{\top}(K+\lambda^{2}I)^{-1}k_{t-1}(x),$
where $K$ is the Gram matrix of the training inputs with $K_{ij}=k(x_{i},x_{j}),\ \forall i,j\in[t-1]$, $k_{t-1}(x)=[k(x,x_{1}),\ldots,k(x,x_{t-1})]^{\top}$ is the covariance vector between the input $x$ and the training inputs, and $y=[y_{1},\ldots,y_{t-1}]^{\top}$ is the vector of observed responses.
The posterior mean $\mu_{t-1}(x)$ and variance $\sigma^{2}_{t-1}(x)$ quantify our posterior belief about the function’s value and uncertainty over $\mathcal{X}$, which we denote compactly as $F_{t-1}=\mathcal{GP}(\mathcal{D}_{t-1})$. While many AFs have been proposed and their utility demonstrated, we focus without loss of generality on the Upper Confidence Bound (UCB) [32], a widely used AF defined as
@@ 2.2 LLM-in-the Loop BO Framework
We start by introducing the general framework. We define the entity running BO as the client. At each iteration $t$, we assume that the client can prompt an LLM agent $\mathcal{A}$, such as a general-purpose model like ChatGPT, to suggest a candidate design to query, denoted $x_{\text{LLM},t}$. This interaction can be implemented using existing approaches [25], or through simple prompt templates tailored to the task at hand [24]. Simultaneously, the client learns the posterior belief via a statistical surrogate conditioned on $\mathcal{D}_{t-1}$ and evaluates $x_{\text{LLM},t}$ accordingly. While our framework does not prescribe a specific surrogate model, we assume without loss of generality that the posterior belief is derived from a $\mathcal{GP}$ model, namely, $F_{t-1}$. Specifically, $F_{t-1}$ contains the information of $\mu_{t-1}(x_{\text{LLM},t})$ and $\sigma_{t-1}^{2}(x_{\text{LLM},t})$, which are used to evaluate $x_{\text{LLM},t}$ with respect to its predicted performance and associated uncertainty. Following this, the client may choose to retain, refine, or reject the $\mathcal{A}$’s suggestion. For now, we describe this decision step only at a high level, as it will be detailed through the three algorithms presented later. This high-level framework is outlined in Algorithm 1.
@@
Input: $\mathcal{D}_{0}$, $T$, LLM Agent $\mathcal{A}$, kernel function $k$.
@@
3:     Compute $x_{\mathcal{GP},t}$ using (1)
@@
Without steps 4–6 in Algorithm 1, this reduces to BO by selecting $x_{t}$ as $x_{\mathcal{GP},t}$, and focusing only on step 4 we recover recent LLM-assisted BO approaches, as in [25, 38]. The added steps aim to guide the sampling decision toward more grounded and theoretically justifiable choices that leverage contextual LLM knowledge along with calibrated $\mathcal{GP}$ surrogates and their uncertainty.
Ultimately, our goal is to establish an upper bound on the cumulative regret for all mechanisms to ensure no regret as $T\to\infty$. We define the instantaneous regret at time $t$ as $r_{t}=f(x^{*})-f(x_{t})$, and the cumulative regret as $R_{T}=\sum_{t=1}^{T}r_{t}$. Our following theoretical developments follow the assumptions below:
@@
Let $\gamma_{t-1}$ denote the maximum information gain after time $t-1$, as defined in Equation (4) of [34]. AF is defined as in (2), where $\beta_{t}$ is defined as
@@
2.3 LLINBO-Transient: Exploration by LLMs then Exploitation by Statistical Surrogates
Perhaps the most natural form of collaboration between an LLM and a BO method is to leverage the LLM’s contextual reasoning early in the process, initially placing greater attention on $x_{\text{LLM},t}$, and gradually transition to the $\mathcal{GP}$’s suggestion $x_{\mathcal{GP},t}$, which is obtained by solving (1) using a $\mathcal{GP}$ surrogate as more data are collected. The $\mathcal{GP}$, with its ability to systematically interpolate observed data and calibrate uncertainty, becomes increasingly reliable for guiding exploitation [28, 13].
@@ 2.3 LLINBO-Transient: Exploration by LLMs then Exploitation by Statistical Surrogates
Theorem 1 (Proof in Appendix A.1).
Suppose that Assumptions 1-2 hold, and LLINBO-Transient is used to generate $x_{t},\ \forall t\in[T]$. Let $p_{t}\in[0,1]$ be chosen such that $1-p_{t}\in\mathcal{O}(1/t)$, then the cumulative regret $R_{T}$ is upper bounded by
@@
The assumption on $p_{t}$ implies that $p_{t}\to 1$ at rate $1-\mathcal{O}\left(\frac{1}{t}\right)$. For example, one may choose $p_{t}=1-\frac{1}{t^{2}}$. With this assumption, the algorithm effectively controls the long-term risk of relying on LLM suggestions throughout the optimization process. Based on this assumption, the proof builds on the high-probability regret analysis framework from [9], leveraging concentration inequalities to control the randomness in selecting evaluation designs under the LLINBO-Transient scheme.

### v2

@@ Abstract
Figure 1: Diagrams of existing methods and the proposed algorithms: LLINBO-Transient, LLINBO-Justify, and LLINBO-Constrained, introduced in Secs. 2.32.5.
@@ 1 Introduction
BO has emerged as a powerful tool for black-box optimization (BBO), providing a principled framework for balancing exploration and exploitation. BO is particularly useful in scenarios where function evaluations are costly, such as in drug discovery (Korovina et al. (2020)), interaction design (Liao et al. (2023)), and hyperparameter tuning (HPT) (Cho et al. (2020)). Starting with an initial dataset, BO employs a surrogate model, most commonly a $\mathcal{GP}$. The $\mathcal{GP}$ is capable of quantifying uncertainty and is used to approximate both the mean and variance of the black-box function. The next query point, hereafter referred to as a design, is then selected by maximizing an acquisition function (AF) that quantifies the potential benefit of evaluating a particular point, thereby strategically balancing exploration and exploitation. BO then augments the dataset with the new design–outcome tuple and proceeds sequentially. The past decade has witnessed many success stories for BO, and its theoretical guarantees have been well established for a range of commonly used AFs (Srinivas et al. (2009); Agrawal & Goyal (2012)). These guarantees are typically regret-based, ensuring that, with high probability, one can asymptotically recover an optimal design.
Recently, the few-shot learning capabilities of LLMs and their ability to generate high-quality outputs from minimal examples have made them attractive tools for optimization tasks (Yang et al. (2024)). In particular, LLMs have shown strong empirical performance over random search (Liu et al. (2024)), largely due to their ability to leverage problem context to fast-track the exploration of promising designs. Intuitively, LLMs act like domain experts, using contextual cues to identify high-quality designs early in the optimization process. At each iteration, different phases of BO, including initial data generation, proposing new designs, and surrogate modeling, are carried out by the LLM through appropriately tailored prompts (Liu et al. (2024); Yang et al. (2024)). These prompts incorporate the current dataset, typically presented as a list of design-response pairs, together with the problem context, enabling the LLM to function as an optimizer. This prompting framework allows LLMs to act as potential agents for BBO without the need for explicit surrogate modeling or large amounts of observed data. We refer to this class of approaches, where LLMs are solely responsible for proposing design candidates and serve as the surrogate model in BO, as LLM-assisted BO.
@@
While recent work on LLM-assisted BO (Liu et al. (2024); Guo et al. (2024); Song et al. (2024); Yang et al. (2024)) has demonstrated promise in generating reasonable query designs, several limitations hinder its broader applicability. Most importantly, LLMs do not provide explicit surrogate modeling or calibrated uncertainty, both of which are essential for principled exploration–exploitation trade-offs. Consequently, although LLMs can accelerate optimization in the early stages, their effectiveness systematically degrades as more data are collected and surrogate models strengthen. As we highlight later, this degradation is a central characteristic that we explicitly model and hedge against in our proposed framework.
Moreover, LLMs remain inherently opaque, making the aforementioned trade-off difficult to interpret or control. This structural opacity, combined with their inability to quantify uncertainty in a principled way, introduces significant risks, particularly in applications where cost or safety is critical, ultimately undermining theoretical tractability and reliability. For instance, in the case of smooth functions, the predictive capability of $\mathcal{GP}$s, in terms of both the predicted mean and variance as measured by generalization bounds, has a known rate of improvement as more data is gathered (Srinivas et al. (2009)). The same result is hard to characterize for LLMs, whose internal mechanisms for interpolating black-box functions are not fully understood and which currently lack calibrated uncertainty estimates.
With this in mind, we propose LLINBO, a framework that combines the contextual reasoning strengths of LLMs with the principled uncertainty quantification offered by statistical surrogates to enable more trustworthy optimization. To operationalize this collaboration, we introduce a general framework grounded in the philosophy of using LLM-suggested designs to sequentially refine and tailor BO. Within this framework, we propose three approaches, which are inspired by recent developments in federated learning, and analyze the theoretical properties of each. Through extensive simulations and a real-world proof-of-concept in 3D printing, we demonstrate the effectiveness and robustness of the proposed methods.
@@
LLMs’ ability to utilize problem context has been actively investigated. Recent work has also demonstrated that LLMs can generalize effectively from limited in-context information (Lampinen et al. (2025); Brown et al. (2020)), making them particularly promising for BBO, where the objective function is unknown and historical observations are limited (Liu et al. (2024)). The use of LLMs for optimization is a growing research direction. An overview of existing LLM-assisted BO can be found in Appendix A.
Based on our best knowledge, the proposed LLINBO is the first hybrid framework that integrates both LLMs and $\mathcal{GP}$s into the BO process to accelerate decision-making. We acknowledge that incorporating external information into BO has been investigated in other settings. For example, in Federated BO (F-BO, Dai et al. (2020); Yue et al. (2025); Chen et al. (2025); Dai et al. (2024)), clients cooperatively perform BO under different sharing schemes. In Human–AI Collaborative BO (HAIC-BO, Hvarfner et al. (2022); Xu et al. (2024); Adachi et al. (2024)), human preferences or belief distributions are incorporated into the BO process. By contrast, the role of LLMs in our framework is fundamentally different from the role of clients in F-BO or humans in HAIC-BO. The few-shot learning ability of LLMs enables the generation of high-quality candidate points in low-data regimes (Liu et al. (2024); Brown et al. (2020)); however, this ability systematically degrades relative to surrogate models as more data accumulate (also demonstrated in our experiments). clients and humans in F-BO and HAIC-BO do not exhibit such properties. This distinction underpins the novelty of our work: LLINBO explicitly models this degradation and introduces principled mechanisms to hedge against LLM unreliability while leveraging their early-stage strengths in tandem with $\mathcal{GP}$s.
[…]
A detailed review of existing F-BO and HAIC-BO is provided in Appendix A; here, we focus on the works that are most directly relevant to the proposed method. In Dai et al. (2020; 2021), Federated Thompson Sampling for BO was introduced, where clients share $\mathcal{GP}$ Random Fourier Features Rahimi & Recht (2007). Each client then selects the next design to query either based on its own features or on those of another randomly chosen client. Alternatively, Chen et al. (2025) proposed a constraint-sharing strategy, where clients resample their surrogates using shared constraints to guide the next evaluation. While our framework differs in its ultimate objective, these principles have directly inspired our hybrid collaboration between LLMs and statistical surrogates.
@@ 2.1 Preliminaries
BO aims to find an optimal design $x^{*}$ that maximizes a black-box function $f$ over a domain $\mathcal{X}$ by sequentially selecting query designs. Given a total budget of $T$ evaluations, the data at iteration $t\in[T]$ is denoted as $\mathcal{D}_{t-1}=\{(x_{i},y_{i})\}_{i=1}^{t-1}$, where $y_{i}=f(x_{i})+\epsilon_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,\lambda^{2})$.
At time $t$, BO selects the next design, denoted by $x_{t}$, to observe by maximizing an AF, $\alpha(x,F_{t-1})$, where $F_{t-1}$ is the posterior belief of $f$ conditioned on $\mathcal{D}_{t-1}$. After selecting $x_{t}$, a noisy observation $y_{t}=f(x_{t})+\epsilon_{t}$ is obtained, and the dataset is updated as $\mathcal{D}_{t}=\mathcal{D}_{t-1}\cup\{(x_{t},y_{t})\}$. This process is then repeated until $T$ is exhausted. The posterior belief is typically modeled using a $\mathcal{GP}$ (Kushner (1964)), which requires a prior mean function $\mu(x)$ (often set to zero) and a kernel function $k(x,x^{\prime})$ encoding the smoothness of the function. This yields a posterior predictive distribution for $f$ given as
@@
with $\mu_{t-1}(x)=k_{t-1}(x)^{\top}(K+\lambda^{2}I)^{-1}y$ and $\sigma^{2}_{t-1}(x)=k(x,x)-k_{t-1}(x)^{\top}(K+\lambda^{2}I)^{-1}k_{t-1}(x),$ where $K$ is the Gram matrix of the training inputs with $K_{ij}=k(x_{i},x_{j}),\ \forall i,j\in[t-1]$, $k_{t-1}(x)=[k(x,x_{1}),\ldots,k(x,x_{t-1})]^{\top}$ being the covariance vector between the input $x$ and the training inputs, and $y=[y_{1},\ldots,y_{t-1}]^{\top}$ is the vector of observed responses.
The posterior mean $\mu_{t-1}(x)$ and variance $\sigma^{2}_{t-1}(x)$ quantify our posterior belief about the function’s value and uncertainty over $\mathcal{X}$, which we denote compactly as $F_{t-1}=\mathcal{GP}(\mathcal{D}_{t-1})$. While many AFs have been proposed and their utility demonstrated, we focus without loss of generality on the Upper Confidence Bound (UCB, Srinivas et al. (2009)), a widely used AF defined as
@@ 2.2 LLM-in-the Loop BO Framework
We start by introducing the general framework and define the entity running BO as the client. At each iteration $t$, we assume that the client can prompt an LLM agent $\mathcal{A}$, such as ChatGPT, to suggest a candidate design to query, denoted $x_{\text{LLM},t}$. This interaction can be implemented using a direct prompt from the client to obtain a query design, or through recently developed approaches and prompt templates tailored to the task at hand (Liu et al. (2024; 2025)).Simultaneously, the client learns the posterior belief via a statistical surrogate conditioned on $\mathcal{D}_{t-1}$ and evaluates $x_{\text{LLM},t}$ accordingly. While our framework does not prescribe a specific surrogate model, we assume without loss of generality that the posterior belief is derived from a $\mathcal{GP}$ model, namely, $F_{t-1}$. Specifically, $F_{t-1}$ contains the information of $\mu_{t-1}(x_{\text{LLM},t})$ and $\sigma_{t-1}^{2}(x_{\text{LLM},t})$, which are used to evaluate $x_{\text{LLM},t}$ with respect to its predicted performance and associated uncertainty. Following this, the client may choose to retain, refine, or reject agent $\mathcal{A}$’s suggestion. For now, we describe this decision step only at a high level in Algorithm 1, as it will be detailed in the three algorithms presented later.
@@
Input: $\mathcal{D}_{0}$, $T$, LLM Agent $\mathcal{A}$, kernel function $k$, AF $\alpha$.
@@
3:  Compute $x_{\mathcal{GP},t}$ by finding the maximizer of $\alpha(x,F_{t-1})$.
@@
Without steps 4–6 in Algorithm 1, this reduces to BO by selecting $x_{t}$ as $x_{\mathcal{GP},t}$, and focusing only on step 4 we recover LLM-assisted BO approaches, as in Liu et al. (2024). The added steps aim to guide the sampling decision toward more grounded and theoretically justifiable choices that leverage contextual LLM knowledge along with calibrated $\mathcal{GP}$ surrogates and their uncertainty.
We define the instantaneous regret at time $t$ as $r_{t}=f(x^{*})-f(x_{t})$, and the cumulative regret as $R_{T}=\sum_{t=1}^{T}r_{t}$. The goal is to establish an upper bound on the $R_{T}$ for all mechanisms to ensure no regret as $T\to\infty$. Our theoretical developments follow the assumptions below:
@@
Let $\gamma_{t-1}$ denote the maximum information gain after time t1t-1, as defined in Equation (4) of Vakili et al. (2021). AF is defined as in (1), where $\beta_{t}$ is defined as
@@
2.3 LLINBO-Transient: Exploration by LLMs then Exploitation by $\mathcal{GP}s$
Perhaps the most natural form of collaboration between an LLM and a BO method is to leverage the LLM’s contextual reasoning early in the process, initially placing greater attention on $x_{\text{LLM},t}$, and gradually transition to the $\mathcal{GP}$’s suggestion $x_{\mathcal{GP},t}$ as more data are collected. The $\mathcal{GP}$, with its ability to systematically interpolate observed data and calibrate uncertainty, becomes increasingly reliable for guiding exploitation (Gramacy (2020)).
@@ 2.3 LLINBO-Transient: Exploration by LLMs then Exploitation by $\mathcal{GP}s$
Theorem 1 (Proof in Appendix B.1).
Suppose that Assumptions 1-2 hold. Let $p_{t}\in[0,1]$ be chosen such that $1-p_{t}\in\mathcal{O}(1/t)$, Then, with probability at least $1-\delta$, $R_{T}$ is upper bounded by
@@
The assumption on $p_{t}$ implies that $p_{t}\to 1$ at rate $1-\mathcal{O}\left(\frac{1}{t}\right)$. For example, one may choose $p_{t}=1-\frac{1}{t^{2}}$. With this assumption, the algorithm effectively controls the long-term risk of relying on LLM suggestions throughout the optimization process. Based on this assumption, we apply the Azuma–Hoeffding inequality introduced in Hoeffding (1963) to upper bound the cumulative regret with high probability, which is a standard technique in the BO literature Dai et al. (2020).